OpenStudy (anonymous):

HELP ME GUYS PLEASE, I'M BEGGING YOU HUHU :( 1. A right circular cylinder has 4in as its diameter and 6in as its height. a. Find the lateral area b. Find the volume of the regular hexagonal prism inscribed inside the cylinder 2. A rectangular solid is 4ft long and 3ft wide a. Find the volume if its diagonal is square root of 41

OpenStudy (anonymous):

lateral surface area= 2 x 3.14 x 4/2 x6

OpenStudy (anonymous):

how about letter b in number 1?

OpenStudy (jdoe0001):

do you know what a "regular hexagonal prism " is?

OpenStudy (anonymous):

i have very cloudy information about it. my teacher is not that good. :(

OpenStudy (anonymous):

also, what's the answer for number 2?

OpenStudy (jdoe0001):

hmm do you know what a right circular cylinder is?

OpenStudy (anonymous):

oh i remember, hexagonal prism has a hexagon as its base. the bases are perpendicular to each other right?

OpenStudy (anonymous):

we are not yet discussing it :(

OpenStudy (jdoe0001):

seems to me you'd need to cover all that first, it doesn't help trying to solve what you haven't dunno what it's

OpenStudy (anonymous):

its an assignment, can you help me answer it please? :(

OpenStudy (anonymous):

also for number 2 please?

OpenStudy (jdoe0001):

a) http://www.mathwords.com/r/right_circular_cylinder.htm

OpenStudy (anonymous):

how about b?

OpenStudy (jdoe0001):

b) http://www.mathwords.com/s/segment_of_a_circle.htm you get the volume of the cylinder then get the 6 segments NOT OCCUPIED by the hexagonal prism and subtract it from the cylinder's volume

OpenStudy (anonymous):

how about number 2?

OpenStudy (jdoe0001):

using the diagonal equation, you can find "h" height you have the length and the width, once you have the height volume = length * width * height

OpenStudy (anonymous):

how do you find the height? what's the formula?

OpenStudy (anonymous):

is it d= square root of(l/^2 +w^2 + h^2? am I right?

OpenStudy (jdoe0001):

A rectangular solid is 4ft long and 3ft wide a. Find the volume if its diagonal is square root of 41 \(\bf diagonal = \sqrt{41}\qquad length=4\qquad width=3\qquad height={\color{red}{ h}}\\ \quad \\ d=\sqrt{l^2+w^2+{\color{red}{ h}}^2}\implies \sqrt{41}=\sqrt{l^2+w^2+{\color{red}{ h}}^2}\implies 41=l^2+w^2+{\color{red}{ h}}^2\)

OpenStudy (jdoe0001):

\(\bf diagonal = \sqrt{41}\qquad length=4\qquad width=3\qquad height={\color{red}{ h}}\\ \quad \\ d=\sqrt{l^2+w^2+{\color{red}{ h}}^2}\implies \sqrt{41}=\sqrt{l^2+w^2+{\color{red}{ h}}^2}\\ \quad \\\implies 41=l^2+w^2+{\color{red}{ h}}^2\)

OpenStudy (anonymous):

oh i see? but i still cannot understand letter b. can you please explain it to me again?

OpenStudy (jdoe0001):

what's the volume of the cylinder?

OpenStudy (jdoe0001):

keeping in mind that cylinder's volume = \(\Large \bf \pi r^2 h\)

OpenStudy (anonymous):

75.40 in^3

OpenStudy (jdoe0001):

ok... .so we know the volume of the cylinder... so let's see an incribed hexagonal a REGULAR hexagonal has 6 EQUAL sides it's a closed polygon so if we divide the circle in 6 EQUAL pieces, each piece will have an angle of 360/6 so we know that the angle for a regular hexagonal is 60 degrees for each side we also know the radius is "2" so  so we want to know the SEGMENT's area that's not occupied by the hexagonal, \(\bf \textit{segment of a circle}=\cfrac{\theta r^2 \pi}{360}\implies \cfrac{{\color{blue}{ 60}} \cdot {\color{red}{ 2}}^2 \pi}{360}\)

OpenStudy (jdoe0001):

the volume of that segment will be the height times the Area, so 6 * Area you have 6 segments, so sum them up, then subtract their volume from the 75.40 from the cylinder, and what's left is the hexagonal's volume

OpenStudy (jdoe0001):

hmm wait a sec... I used .. .dohh the wrong ... equation for the segment, I got the SECTOR instead, lemme rewrite that

OpenStudy (jdoe0001):

so 60 degrees and radius =2 \(\bf \cfrac{r^2}{3}\left(\cfrac{\pi \theta}{180}-sin(\theta)\right)\implies \cfrac{{\color{blue}{ 2}}^2}{3}\left(\cfrac{\pi \cdot {\color{red}{ 60}}}{180}-sin({\color{red}{ 60^o}})\right)\)

OpenStudy (anonymous):

so what's your asnwer in lettter b?

OpenStudy (jdoe0001):

the volume of that segment will be the height times the Area, so 6 * Area you have 6 segments, so sum them up, then subtract their volume from the 75.40 from the cylinder, and what's left is the hexagonal's volume

OpenStudy (anonymous):

is your asnwer 66.7?

OpenStudy (jdoe0001):

hmmm

OpenStudy (jdoe0001):

I just notice a typo ark

OpenStudy (jdoe0001):

my bad

OpenStudy (jdoe0001):

\(\bf \cfrac{r^2}{2}\left(\cfrac{\pi \theta}{180}-sin(\theta)\right)\implies \cfrac{{\color{blue}{ 2}}^2}{2}\left(\cfrac{\pi \cdot {\color{red}{ 60}}}{180}-sin({\color{red}{ 60^o}})\right)\)

OpenStudy (jdoe0001):

\(\bf \cfrac{r^2}{2}\left(\cfrac{\pi \theta}{180}-sin(\theta)\right)\implies \cfrac{{\color{blue}{ 2}}^2}{2}\left(\cfrac{\pi \cdot {\color{red}{ 60}}}{180}-sin({\color{red}{ 60^o}})\right)\implies 2\left(\cfrac{\pi}{3}-sin(60^o)\right)\\ \quad \\ 2\left(1.047-\frac{\sqrt{3}}{2}\right)\approx 0.3623\\ \quad \\ \textit{volume of the segment will be } 0.3623 \cdot 6\implies 2.1741\\ \quad \\ \textit{6 segments, so }2.1741\cdot 6 = 13.0444\\ \quad \\ {\color{red}{ 75.4}}-13.044 \approx 62.3556\)

OpenStudy (anonymous):

THANK YOU SO MUCH! :D

OpenStudy (jdoe0001):

yw